# T-Shirt: Token-Selective Hierarchical Data Selection for Instruction Tuning

### v1

@@ 3.2 Hierarchical Selection Based on Local Neighborhood
[…]
@@
[…]
The sensitivity of IFD and S-IFD scores to small input shifts motivates us to rethink these scores. Commonly, they regard training samples as high- or low-quality only based on whether their quality scores (e.g., IFD or S-IFD) exceed a given threshold [16, 17], which has several drawbacks as we illustrate in Figure 4. For example, sample $B$ may not be a reliable high-quality sample because many nearby points have low scores (essentially easy). This suggests that high average quality among neighbors is desirable for a sample. However, average quality alone might be insufficient. For instance, sample $C$ lies in a region with high variance, indicating instability. In contrast, the neighbors of sample $A$ exhibit both high average scores and low local variance, making it a more robust candidate. Thus, we propose selecting samples based not only on high local average quality but also on low variance within their neighborhood.
Our Approach  Motivated by our observation that local properties, including the average and variance of a sample’s neighbors’ S-IFD scores, are valuable for selection, we propose to use these measures as an additional dimension for evaluating data quality. To estimate these properties for a training sample $(x,y)$, we first generate its neighbors. Rather than using complex lexical perturbations such as paraphrasing or word substitution [30, 31], we opt for a simpler approach: injecting random noise into the token embeddings. We add uniformly distributed noise to each token embedding. For each token, we sample a noise vector with entries uniformly drawn i.i.d. from $[-1,1]$, then scale it by $\epsilon=\alpha/\sqrt{(L+T)d}$, where $\alpha$ is a hyperparameter, $L$ and $T$ are the lengths of the instruction and response, respectively, and $d$ is the embedding dimension. This scaling results in perturbed samples whose expected $\ell_{2}$-norm distance from the original sample is $\alpha/\sqrt{3}$ [33]. We denote the perturbed embeddings of $(x,y)$ as $(x+\delta_{x}^{(i)},y+\delta_{y}^{(i)})$, and compute the S-IFD score for each perturbed sample. We repeat this process $M$ times to estimate the average and variance of the S-IFD scores: [^1] For simplicity, we use $\hat{\mu}(x,y)$ to denote $\mu(\text{S-IFD}_{k}(x,y))$, and $\hat{\sigma}^{2}(x,y)$ to denote $\sigma^{2}(\text{S-IFD}_{k}(x,y))$.
@@
After computing these local properties, we apply a hierarchical selection strategy to identify high-quality training samples. Given a budget $b$, we first select $\gamma b$ samples whose neighbors have the highest average S-IFD scores, where $\gamma>1$ denotes an upsampling factor. From these, we further select the final $b$ samples with the lowest variance in S-IFD scores among their neighbors. We outline our data selection method in Algorithm 1 and visualize it in Figure 1. In practice, to improve efficiency, we batch the noisy embeddings instead of using a naive for-loop.
@@
[…]
      Compute $\hat{\mu}(x,y)$ and $\hat{\sigma}^{2}(x,y)$ via Equation 4
@@ 4.1 Experimental Setup
Instruction Tuning Datasets  We conduct data selection on two widely used instruction-tuning datasets of different initial qualities and scales: Alpaca-GPT-4 [11] and Magpie [10]. Following prior works [15, 16, 17], we use various data selection methods to select 5% of the 52k samples in Alpaca-GPT-4 and approximately 3.3% (10k samples) of the 300k-sample Magpie dataset. These sampling ratios have been shown to reach an effective balance between model performance and data efficiency [15, 16, 17]. We provide additional details on these two datasets in Section C.1.
@@
Baselines  On Alpaca-GPT-4, we compare our method against the following baselines: (1) Full, which uses the entire dataset for instruction tuning; (2) Random, which randomly selects data from the full training set; (3) Longest [23], which selects samples with the longest response lengths; (4) Deita [14], which first trains LLM scorers on ChatGPT-labeled data to assess sample quality and complexity, then selects data based on quality, complexity, and diversity; (5) DS2 [15], which first uses GPT-4o-mini to score training samples, then learns a score transition matrix to correct potential scoring errors, and finally selects data based on both quality and diversity; (6) IFD [17], which uses GPT-2-computed IFD scores [16] to select high-quality samples. On Magpie, due to limited resources, we evaluate only the best-performing baselines identified on Alpaca-GPT-4. More details about the baselines are provided in Section C.2.
@@
Benchmarks  We evaluate the performance of the trained LLMs across eight diverse benchmarks. These include six standardized benchmarks from the OpenLLM Leaderboard [34, 35]: ARC-Challenge [36], HellaSwag [37], MMLU [38], TruthfulQA [39], BBH [40], and GSM8k [41]. For these benchmarks, we use the LM-Evaluation-Harness [42] and report their default evaluation metrics. We denote their average score as $\mu_{\textsc{open}}$. In addition, we assess the instruction-tuned LLMs on two LLM-as-a-Judge benchmarks: Arena-Hard [43] and AlpacaEval 2.0 [44]. Following the recommendations of the respective benchmark creators, we report the style-controlled win rate [43] for Arena-Hard and the length-controlled win rate [44] for AlpacaEval 2.0. The average score across these two is denoted as $\mu_{\textsc{llm}}$. Finally, we report the overall average across all eight benchmarks as $\mu_{\textsc{all}}$. Additional details about the benchmarks are provided in Section C.3.
@@
Implementation Details of T-Shirt  To implement our method, T-Shirt, we use GPT-2 to compute S-IFD scores. For the hyperparameters in Algorithm 1, we set the token selection ratio $k\%$ to either 50% or 75%, the oversampling factor to $\gamma=2$, the base noise scale to $\alpha=5$ following NeftTune [33], and the number of perturbations to $M=30$. We include more implementation details in Section C.4.
@@
Instruction Tuning Details  We use selected data to fine-tune two open-source LLMs, Llama-3.1-8B [3] and Qwen-2.5-7B [4]. Following the same training settings as prior works [16, 17, 45], we use a learning rate of 2e-5 and train for 3 epochs. For data selected from Magpie, we instead follow the setup in Magpie [10] and train for 2 epochs. Additional training details are provided in Section C.5.
@@ 4.2 Main Experimental Results
Table 2: Performance comparison between our method, T-Shirt, and other baseline methods on Magpie.  We use each method to select 10k samples from Magpie to instruction-tune Qwen-2.5-7B. We use abbreviated benchmark names as shown in Table 1. The best results are shown in bold, and the second-best are underlined. Our method achieves the best performance on seven out of eight benchmarks and has the highest $\mu_{\textsc{all}}$.
@@
T-Shirt outperforms all baselines.  We evaluate the performance of our method, T-Shirt, and six baseline methods on the Alpaca-GPT-4 dataset using eight benchmarks, with results presented in Table 1. Across two base models, T-Shirt with different token-selection ratios ($k\%$) consistently outperforms all baselines in terms of $\mu_{\textsc{all}}$. Our data selection method achieves the highest performance on both the OpenLLM Leaderboard and LLM-as-a-Judge benchmarks. Specifically, T-Shirt achieves the highest $\mu_{\textsc{open}}$ and $\mu_{\textsc{llm}}$ scores on Qwen-2.5-7B, as well as the highest $\mu_{\textsc{open}}$ and competitive $\mu_{\textsc{llm}}$ performance on Llama-3.1-8B. Compared to the strongest baseline, IFD, T-Shirt delivers superior results on six out of eight benchmarks and performs comparably on the remaining two. These results demonstrate that our method significantly enhances the overall quality of selected data. Notably, selecting only 5% of the data, T-Shirt improves $\mu_{\textsc{all}}$ by 5.10 to 5.48 points compared to instruction tuning using the full dataset.
@@
T-Shirt is effective at selecting data across datasets of varying initial qualities and scales.  We further use T-Shirt, and the two best-performing baselines from Table 1 to select 10k samples from Magpie, a recent state-of-the-art instruction tuning dataset. Magpie is approximately 5.8×5.8$\times$5.8 × larger than Alpaca-GPT-4 and has undergone multi-stage filtering to select 300k samples from an initial pool of one million, leading to higher initial quality. We use the selected data to fine-tune Qwen-2.5-7B and report results in Table 2. Across eight benchmarks, T-Shirt outperforms the baselines on seven of them and achieves the best overall scores in $\mu_{\textsc{open}}$, $\mu_{\textsc{llm}}$, and $\mu_{\textsc{all}}$. These results demonstrate that T-Shirt scales effectively and remains beneficial even on large datasets with high initial quality.
T-Shirt is both cost-effective and computationally efficient.  When having strong performance, T-Shirt relies only on a lightweight GPT-2 model to compute S-IFD scores, avoiding the need for API credits required by Deita and DS2. This makes our method both accessible and efficient. We present runtime comparisons in Table 4. On a single NVIDIA A6000 GPU, T-Shirt runs 2.7 to 3.7$\times$ faster than Deita and DS2. Although it is slower than Longest and IFD, its total runtime remains about 40 minutes, which is an acceptable trade-off given its superior overall performance. Moreover, we can further improve its efficiency by reducing the number of perturbations $M$ without hurting the performance (see ablation study on $M$ in Table 4).
@@
Proprietary LLMs are not (all) you need for data selection.  As shown in Table 1, the baselines Longest, IFD, and our method T-Shirt outperform both Deita and DS2. Longest uses a single heuristic: response length. IFD and T-Shirt assess data quality using only a weak model, GPT-2. In contrast, Deita and DS2 rely on expensive proprietary LLMs and employ more complex data selection strategies. Moreover, data selected by Deita and DS2 might contain intrinsic biases, because both of them show consistently lower performance on GSM8k, and Deita also underperforms on TruthfulQA, regardless of the base LLM used for instruction tuning.
@@ 4.3 Ablation Studies and Analysis
In this section, we conduct ablation studies using the Alpaca-GPT-4 dataset and OpenLLM Leaderboard benchmarks. Additional details are provided in Appendix D.
@@
Ablation on S-IFD and Hierarchical Selection  In Table 4, we evaluate the contributions of S-IFD and hierarchical selection. The results show that, compared to the baseline IFD, both components significantly improve $\mu_{\textsc{open}}$ of instruction-tuned LLMs, highlighting the importance of token-level informativeness and local quality consistency in data selection. Combining both factors, our method achieves the best overall performance.
@@
Ablation on Neighbor Variance in Hierarchical Selection  As shown in Figure 5 in Figure 5, $\mu_{\textsc{open}}$ drops significantly, by 2.29 to 4.86 points, when hierarchical selection chooses samples whose neighbors have high variance in S-IFD. This further highlights the importance of choosing data whose neighbors display consistently high quality.
@@
Ablation on Token Selection RatioFigure 6(a) in Figure 6 shows how $\mu_{\textsc{open}}$ of LLMs changes with different token selection ratios $k\%$. For Llama-3.1-8B, the optimal ratio is 75%, while for Qwen-2.5-7B, it is 50%. These results support our observation in Section 3.1 that over 20% of the tokens in the instruction tuning dataset are not informative for effective data selection.
@@
Ablation on Number of PerturbationsFigure 6(b) in Section D.3 illustrates the impact of varying the number of perturbations $M$ used in hierarchical selection. We observe that $\mu_{\textsc{open}}$ for Llama-3.1-8B reaches its highest value at $M=10$, while $\mu_{\textsc{open}}$ for Qwen-2.5-7B peaks at $M=20$. Both models perform well at $M=30$. This suggests that reducing $M$ to lower values can improve the efficiency of our method without compromising performance.

### v2

@@ 3.2 Hierarchical Selection Based on Local Neighborhood
Figure 3: Sensitivity of IFD and $\text{S-IFD}_{75}$ scores to a semantics-preserving word substitution.Top: An instruction-response pair with a blank, filled with either “average” or “mean”. Bottom: Plot of $|\Delta_{t}|$ values for each variant.
@@
We note that LLMs can be highly sensitive to subtle input changes, even those that preserve semantics. As a result, for an instruction-response pair $(x,y)$, the computation of $\Delta_{t}$ for token $y_{t}$ may vary significantly with minor changes in the instruction $x$ or the preceding response tokens $y_{<t}$. These variations can accumulate, leading to substantial shifts in IFD and S-IFD scores. Even small perturbations to $(x,y)$ can lead to large fluctuations in IFD and S-IFD scores, as we illustrate through an example from the Alpaca-GPT-4 dataset in Figure˜4. Here, the instruction asks for the average of four integers. When the word “average” is replaced with the synonym “mean”, both IFD and $\text{S-IFD}_{75}$ scores drop sharply. This highlights the model’s sensitivity to superficial lexical variations and suggests that high scores may sometimes capture sensitivity to spurious surface features rather than genuine semantic quality.
The sensitivity of IFD and S-IFD scores to small input shifts motivates us to rethink these scores. Commonly, they regard training samples as high- or low-quality only based on whether their quality scores (e.g., IFD or S-IFD) exceed a given threshold [li2024ifd, li2024superfiltering], which has several drawbacks as we illustrate in Figure˜4. For example, sample $B$ may not be a reliable high-quality sample because many nearby points have low scores (essentially easy). This suggests that high average quality among neighbors is desirable for a sample. However, average quality alone might be insufficient. For instance, sample $C$ lies in a region with high variance, indicating instability. In contrast, the neighbors of sample $A$ exhibit both high average scores and low local variance, making it a more robust candidate. Thus, we propose selecting samples based not only on high local average quality but also on low variance within their neighborhood.
Our Approach  Motivated by our observation that local properties, including the average and variance of a sample’s neighbors’ S-IFD scores, are valuable for selection, we propose to use these measures as an additional dimension for evaluating data quality. To estimate these properties for a training sample $(x,y)$, we first generate its neighbors. Rather than using complex lexical perturbations such as paraphrasing or word substitution [zang2020word, wang2021adversarial], we opt for a simpler approach: injecting random noise into the token embeddings. We add uniformly distributed noise to each token embedding. For each token, we sample a noise vector with entries uniformly drawn i.i.d. from [1,1][-1,1], then scale it by $\epsilon=\alpha/\sqrt{(L+T)d}$, where $\alpha$ is a hyperparameter, $L$ and $T$ are the lengths of the instruction and response, respectively, and $d$ is the embedding dimension. This scaling results in perturbed samples whose expected $\ell_{2}$-norm distance from the original sample is $\alpha/\sqrt{3}$ [jain2024neftune]. We denote the perturbed embeddings of $(x,y)$ as $(x+\delta_{x}^{(i)},y+\delta_{y}^{(i)})$, and compute the S-IFD score for each perturbed sample. We repeat this process $M$ times to estimate the average and variance of the S-IFD scores: [^1] For simplicity, we use $\hat{\mu}(x,y)$ to denote $\mu(\text{S-IFD}_{k}(x,y))$, and $\hat{\sigma}^{2}(x,y)$ to denote $\sigma^{2}(\text{S-IFD}_{k}(x,y))$.
@@
After computing these local properties, we apply a hierarchical selection strategy to identify high-quality training samples. Given a budget $b$, we first select $\gamma b$ samples whose neighbors have the highest average S-IFD scores, where $\gamma>1$ denotes an upsampling factor. From these, we further select the final $b$ samples with the lowest variance in S-IFD scores among their neighbors. We outline our data selection method in Algorithm˜1 and visualize it in Figure˜1. In practice, to improve efficiency, we batch the noisy embeddings instead of using a naive for-loop.
@@
      Compute $\text{S-IFD}_{k}(x^{\prime},y^{\prime})$ via Equation˜3
  Compute $\hat{\mu}(x,y)$ and $\hat{\sigma}^{2}(x,y)$ via Equation˜4
@@ 4.1 Experimental Setup
Instruction Tuning Datasets  We conduct data selection on two widely used instruction-tuning datasets of different initial qualities and scales: Alpaca-GPT-4 [peng2023instructiontuninggpt4] and Magpie [xu2025magpie]. Following prior works [pang2025ds2, li2024ifd, li2024superfiltering], we use various data selection methods to select 5% of the 52k samples in Alpaca-GPT-4 and approximately 3.3% (10k samples) of the 300k-sample Magpie dataset. These sampling ratios have been shown to reach an effective balance between model performance and data efficiency [pang2025ds2, li2024ifd, li2024superfiltering]. We provide additional details on these two datasets in Section˜B.1.
@@
Baselines  On Alpaca-GPT-4, we compare our method against the following baselines: (1) Full, which uses the entire dataset for instruction tuning; (2) Random, which randomly selects data from the full training set; (3) Longest [zhao2024long], which selects samples with the longest response lengths; (4) Deita [liu2024deita], which first trains LLM scorers on ChatGPT-labeled data to assess sample quality and complexity, then selects data based on quality, complexity, and diversity; (5) DS2 [pang2025ds2], which first uses GPT-4o-mini to score training samples, then learns a score transition matrix to correct potential scoring errors, and finally selects data based on both quality and diversity; (6) IFD [li2024superfiltering], which uses GPT-2-computed IFD scores [li2024ifd] to select high-quality samples. On Magpie, due to limited resources, we evaluate only the best-performing baselines identified on Alpaca-GPT-4. More details about the baselines are provided in Section˜B.2.
@@
Benchmarks  We evaluate the performance of the trained LLMs across eight diverse benchmarks. These include six standardized benchmarks from the OpenLLM Leaderboard [open-llm-leaderboard-v1, open-llm-leaderboard-v2]: ARC-Challenge [allenai:arc], HellaSwag [zellers2019hellaswag], MMLU [hendrycks2021mmlu], TruthfulQA [lin2022truthfulqa], BBH [suzgun2022bbh], and GSM8k [cobbe2021gsm8k]. For these benchmarks, we use the LM-Evaluation-Harness [eval-harness] and report their default evaluation metrics. We denote their average score as $\mu_{\textsc{open}}$. In addition, we assess the instruction-tuned LLMs on two LLM-as-a-Judge benchmarks: Arena-Hard [li2024arenahard] and AlpacaEval 2.0 [dubois2024alpacaeval2]. Following the recommendations of the respective benchmark creators, we report the style-controlled win rate [li2024arenahard] for Arena-Hard and the length-controlled win rate [dubois2024alpacaeval2] for AlpacaEval 2.0. The average score across these two is denoted as $\mu_{\textsc{llm}}$. Finally, we report the overall average across all eight benchmarks as $\mu_{\textsc{all}}$. Additional details about the benchmarks are provided in Section˜B.3.
@@
Implementation Details of T-Shirt  To implement our method, T-Shirt, we use GPT-2 to compute S-IFD scores. For the hyperparameters in Algorithm˜1, we set the token selection ratio $k\%$ to either 50% or 75%, the oversampling factor to $\gamma=2$, the base noise scale to $\alpha=5$ following NeftTune [jain2024neftune], and the number of perturbations to $M=30$. We include more implementation details in Section˜B.4.
@@
Instruction Tuning Details  We use selected data to fine-tune two open-source LLMs, Llama-3.1-8B [grattafiori2024llama3herdmodels] and Qwen-2.5-7B [qwen2025qwen25technicalreport]. Following the same training settings as prior works [li2024ifd, li2024superfiltering, Taori2023alpaca], we use a learning rate of 2e-5 and train for 3 epochs. For data selected from Magpie, we instead follow the setup in Magpie [xu2025magpie] and train for 2 epochs. Additional training details are provided in Section˜B.5.
@@ 4.2 Main Experimental Results
In this section, we present the experimental results to demonstrate the effectiveness of T-Shirt. Additional results are provided in Appendix˜C.
Table 2: Performance comparison between our method, T-Shirt, and other baseline methods on Magpie.  We use each method to select 10k samples from Magpie to instruction-tune Qwen-2.5-7B. We use abbreviated benchmark names as shown in Table˜1. The best results are shown in bold, and the second-best are underlined. Our method achieves the best performance on seven out of eight benchmarks and has the highest $\mu_{\textsc{all}}$.
@@
T-Shirt outperforms all baselines.  We evaluate the performance of our method, T-Shirt, and six baseline methods on the Alpaca-GPT-4 dataset using eight benchmarks, with results presented in Table˜1. Across two base models, T-Shirt with different token-selection ratios ($k\%$) consistently outperforms all baselines in terms of $\mu_{\textsc{all}}$. Our data selection method achieves the highest performance on both the OpenLLM Leaderboard and LLM-as-a-Judge benchmarks. Specifically, T-Shirt achieves the highest $\mu_{\textsc{open}}$ and $\mu_{\textsc{llm}}$ scores on Qwen-2.5-7B, as well as the highest $\mu_{\textsc{open}}$ and competitive $\mu_{\textsc{llm}}$ performance on Llama-3.1-8B. Compared to the strongest baseline, IFD, T-Shirt delivers superior results on six out of eight benchmarks and performs comparably on the remaining two. These results demonstrate that our method significantly enhances the overall quality of selected data. Notably, selecting only 5% of the data, T-Shirt improves $\mu_{\textsc{all}}$ by 5.10 to 5.48 points compared to instruction tuning using the full dataset.
@@
T-Shirt is effective at selecting data across datasets of varying initial qualities and scales.  We further use T-Shirt, and the two best-performing baselines from Table˜1 to select 10k samples from Magpie, a recent state-of-the-art instruction tuning dataset. Magpie is approximately $5.8\times$ larger than Alpaca-GPT-4 and has undergone multi-stage filtering to select 300k samples from an initial pool of one million, leading to higher initial quality. We use the selected data to fine-tune Qwen-2.5-7B and report results in Table˜2. Across eight benchmarks, T-Shirt outperforms the baselines on seven of them and achieves the best overall scores in $\mu_{\textsc{open}}$, $\mu_{\textsc{llm}}$, and $\mu_{\textsc{all}}$. These results demonstrate that T-Shirt scales effectively and remains beneficial even on large datasets with high initial quality. We also fine-tune Qwen-2.5-14B using T-Shirt-selected samples from Magpie and show detailed results in Table˜5 and Section˜C.2.
[…]
@@
Proprietary LLMs are not (all) you need for data selection.  As shown in Table˜1, the baselines Longest, IFD, and our method T-Shirt outperform both Deita and DS2. Longest uses a single heuristic: response length. IFD and T-Shirt assess data quality using only a weak model, GPT-2. In contrast, Deita and DS2 rely on expensive proprietary LLMs and employ more complex data selection strategies. Moreover, data selected by Deita and DS2 might contain intrinsic biases, because both of them show consistently lower performance on GSM8k, and Deita also underperforms on TruthfulQA, regardless of the base LLM used for instruction tuning.
@@ 4.3 Ablation Studies and Analysis
In this section, we conduct ablation studies using the Alpaca-GPT-4 dataset and OpenLLM Leaderboard benchmarks. Additional details are provided in Table˜7.
@@
Ablation on S-IFD and Hierarchical Selection  In Table˜4, we evaluate the contributions of S-IFD and hierarchical selection. The results show that, compared to the baseline IFD, both components significantly improve $\mu_{\textsc{open}}$ of instruction-tuned LLMs, highlighting the importance of token-level informativeness and local quality consistency in data selection. Combining both factors, our method achieves the best overall performance.
@@
Ablation on Neighbor Variance in Hierarchical Selection  As shown in Figure˜6 in Section˜D.1, $\mu_{\textsc{open}}$ drops significantly, by 2.29 to 4.86 points, when hierarchical selection chooses samples whose neighbors have high variance in S-IFD. This further highlights the importance of choosing data whose neighbors display consistently high quality.
@@
Ablation on Token Selection RatioFigure˜7(a) in Section˜D.2 shows how $\mu_{\textsc{open}}$ of LLMs changes with different token selection ratios $k\%$. For Llama-3.1-8B, the optimal ratio is 75%, while for Qwen-2.5-7B, it is 50%. These results support our observation in Section˜3.1 that over 20% of the tokens in the instruction tuning dataset are not informative for effective data selection.
Ablation on Token Weights  In Equation˜3, we assign a weight ($w_{t}$) of 1 to informative tokens and 0 to uninformative ones, effectively removing the latter from the quality score computation. In Table˜7 (Section˜D.3), we compare this binary weighting scheme with alternatives that upweight informative tokens instead of fully discarding uninformative ones. Specifically, we experiment with informative-to-uninformative token weight ratios of 1.5:1 and 2:1. Our results show that the binary weighting scheme consistently outperforms these softer alternatives. This suggests that completely ignoring uninformative tokens yields a cleaner and more reliable signal for computing quality scores.
@@
Ablation on Number of PerturbationsFigure˜7(b) in Section˜D.4 illustrates the impact of varying the number of perturbations $M$ used in hierarchical selection. We observe that $\mu_{\textsc{open}}$ for Llama-3.1-8B reaches its highest value at $M=10$, while $\mu_{\textsc{open}}$ for Qwen-2.5-7B peaks at $M=20$. Both models perform well at $M=30$. This suggests that reducing $M$ to lower values can improve the efficiency of our method without compromising performance.
5 Discussion
5.1 Token Embedding Perturbations vs. Lexical Perturbations
In the hierarchical selection process of T-Shirt, we choose to generate neighbors of each sample using continuous token embedding perturbations rather than lexical perturbations. This decision is motivated by several challenges associated with lexical perturbations.
Utility  Rule-based lexical perturbations often struggle to preserve semantic equivalence and fluency. As a result, prior studies [jin2020isbert, wang2021adversarial] typically rely on costly human evaluations to assess the quality and utility of perturbed examples.
Efficiency  Generating lexical variants by prompting large language models through APIs is both computationally and financially expensive. For example, perturbing each of the 52k Alpaca-GPT4 samples 30 times would require over 1.5 million API calls.
5.2 Limitations and Future Work
While our method shows promising results in data selection for instruction tuning, it has several limitations, which we outline along with directions for future work.
Model Scale  Due to computational constraints, our experiments are limited to models in the 7B–14B parameter range, as is common in several published papers in instruction tuning [xu2025magpie, pang2025ds2, pang2025tokencleaning]. We will apply our data selection method to larger models in future work.
Data Security Considerations  Our approach focuses solely on identifying high-quality subsets of instruction tuning datasets rather than security concerns. Future work could extend our method to incorporate data selection strategies that explicitly focus on security in LLMs.
Theoretical Foundations  Our hierarchical selection method is inspired by research in adversarial machine learning [cohen2019certified, raghunathan2018semidefinite], robustness measures [hamman2023robust, pmlr-v244-han24a], and LLM robustness [hamman2025quantifying, jain2024neftune]. Empirical results support its effectiveness. However, a deeper theoretical understanding of our approach for data selection is still lacking. In future work, we aim to establish a stronger theoretical foundation to better explain our method and potentially even improve its performance.